OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2.Describe how you would graph this line using the slope-intercept method. Be s

OpenStudy (anonymous):

1. slope intercept form is Y = mx + b so we are going to rearrange the given equation into that form given: 2x + 3y = 1200 subtract 2x from both sides and then divide by 3 3y = -2x + 1200 y = -2/3x + 400

OpenStudy (anonymous):

Did you follow that?

OpenStudy (anonymous):

what is the slope? M y intercept? B

OpenStudy (anonymous):

i think the slope for m is 400

OpenStudy (anonymous):

Look at Y = mx + b m =slope B = y intercept so the slope is actually -2/3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now what is the y intercept?

OpenStudy (anonymous):

and what point (for graphing purposes) does that represent?

OpenStudy (anonymous):

400

OpenStudy (anonymous):

im sorry im not good at math

OpenStudy (anonymous):

nP

OpenStudy (anonymous):

At least you are asking for help :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The point that coincides with the y intercept is (0, B) so that is (0,400) in this problem

OpenStudy (anonymous):

So to answer the last question...always start with y intercept point and then use the slope to RISE (numerator) and RUN RIGHT (denominator). Note when the slope is NEGATIVE RUN to the LEFT.... Draw a line to connect the points and there you have your graph.

OpenStudy (anonymous):

so for question 1 Y = mx + b that was the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you could you help me with one more and how do i give you a medal

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

You r welcome...sure post new one I am here

OpenStudy (anonymous):

3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. • You may graph your equation by hand on a piece of paper and scan your work. • You may graph your equation using graphic technology that can be found in the Course Information area. 5. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences.

OpenStudy (anonymous):

Are you using the same information?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i hate math but is neccesary lol

OpenStudy (anonymous):

3.Solve the equation for y and then replace y by f(x) Since we solved for y = mx + b ...just rewrite to fx) = mx + b

OpenStudy (anonymous):

The graph of the function represents the graph of the line...

OpenStudy (anonymous):

4 we discussed that earlier...just graph the line by using 2 points. One point is the y intercept ordered pair and the other is where the rise/run stops when you apply slope.

OpenStudy (anonymous):

soo (0,400 that is the y

OpenStudy (anonymous):

5. instead of profit being 1200 it is 1500 so equation is 2x + 3y = 1500

OpenStudy (anonymous):

yes, 0,400 is the ordered pair for the y - intercept

OpenStudy (anonymous):

5 cont...from the new equation - rewrite into slope intercept form. compare this one with the one we did in 1...what is different, what is same. look at slope and y intercept.

OpenStudy (anonymous):

wait i dont get it sorry

OpenStudy (anonymous):

Don't get which part?

OpenStudy (anonymous):

the things i need to graph

OpenStudy (anonymous):

plot point (o,b) make it easy - pick a point for x...say x = 30 plug in value for x and what do you get for Y? that is point #2 connect those two points

OpenStudy (anonymous):

so 0,30 that what i plug in

OpenStudy (anonymous):

ordered pairs are x,y so if x = 30 ; what is y?

OpenStudy (anonymous):

plug in x value to determine Y

OpenStudy (anonymous):

im sorry put i dont understand how to plug the x to the y

OpenStudy (anonymous):

y = -2/3x +400 y = -2/3 (30) + 400 y = -20 + 400 y = 380 so the 2nd point is 30,380

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

so is there away that you could draw the graph so i could see it it kind makes sense

OpenStudy (anonymous):

The points are too big. It is actually just a line

OpenStudy (anonymous):